OpenStudy (anonymous):

I have the function \[f(x)=log(x)\] and I have found \[f' = 1/x\] \[f'' = -1/x^2\] \[f''' = 1*2/x^3\] and then i find the general: \[f^n = (-1)^n-1 * n-1!/x^n\] But I need to proof this, can any one help me with this?

geerky42 (geerky42):

To make general case look little nicer: \(f^{(n)}(x) = \dfrac{(-1)^{n-1}(n-1)!}{x^n}\quad\quad n\in\mathbb N\) Not sure how to prove. Can you help? @Loser66 @SithsAndGiggles

OpenStudy (loser66):

I think we can apply induction here.

OpenStudy (badhi):

lets see if its true for n = 1, \[f^{(1)}(x) = \frac{(-1)^{1-1}(1-1)!}{x^1} = \frac{1\times 0!}{x} = \frac{1}{x}\] so it is true for \(n=1\) assume its true for \(n=p\), then \[f^{(p)} = \frac{(-1)^{p-1}(p-1)!}{x^p}\] if we differentiate above wrt x \[\begin{align} \frac{df^{(p)}}{dx} &= \frac{(-1)^{p-1}(p-1)!}{x^p}\\ f^{(p+1)}&=(-1)^{p-1}(p-1)!\left[\frac{dx^{-p}}{dx}\right]\\ &=(-1)^{p-1}(p-1)!\left[(-p)x^{-(p+1)}\right] \end{align}\] See if you can simplify it to following \[f^{(p)} = \frac{(-1)^{p}(p)!}{x^{p+1}}\] if you can, you have proven that if the expression is true for \(n=p\) it is also true for \(n=p+1\) since at the begining we have proven that the expression is true for \(n=1\) we can conclude that, \[f^{(n)} = \frac{(-1)^{n-1}(n-1)!}{x^n}\] is true for \(\forall x \in \mathbb N\)

OpenStudy (badhi):

oops @Loser66

OpenStudy (loser66):

oh, yeah, you got it. hehehe

OpenStudy (loser66):

just take derivative of f^(n) to get the answer. it is not hard, right?

OpenStudy (badhi):

yeah, thats the main part.. but the process should be followed as well :)